OpenStudy (anonymous):

A line passes through the point (0, –4) and has a slope of 3/2. What is the equation of the line? A. 3x-2y=8 B. 6x+4y=-8 C. 2x+4y=-1 D. 3x-2y=8

OpenStudy (anonymous):

y-y1=m(x=x1) m is the slope =3/2 (x1,y1) is the point which is (0,4) y-4=3/2(x-0)

OpenStudy (anonymous):

y=3/2+4

OpenStudy (anonymous):

multiply by 2 2y=3x+8